A body is tied to one end of a string and whirled in a vertical plane in a circle. calculate the minimum velocity at highest and lowest point of the circle for loopinf the vertical loop.given the radius of the circle is r

Grade:11

3 Answers

Adarsh
6 years ago
Dear Yash,
it can be calculated by conserving energy ….
1/2m.v2=m.g.2r =>  v = 4gr 
Thanks and Regards
Adarsh
Chandan
6 years ago
at highest point:
mg=mv2/r
v=(gr)1/2
at lowest point:
½ mgr+mg2r=1/2mv2
v=(4gr)1/2
hope it helps
please approve if it does
 
Upendra Yadav
6 years ago
2at lowest point:...½ mgr+mg2r=1/2mv½.v=(gr)/r2 mg=mv
v=(4gr)½
REGARDS WITH 
U[ENDRA YADAV